Two waves travelling in the same direction are given by: y1 = 0.2 sin⁡(2πx-20t+φ) and y2 = 0.2 sin⁡(2πx-20t), where x and y are in meters and t is in seconds. If the two waves start at the same moment, then the path difference, ∆x, corresponding to a fully destructive interference is:

2/3 m


3/4 m


1/3 m


1/2 m


4/3 m


1
Expert's answer
2021-05-20T10:09:44-0400

"y_1=0.2\sin(2\pi x-20t+\phi)"

"y_2=0.2\sin(2\pi x-20t)"

"y=y_1+y_2"

"y=\bigg(0.2\cos\dfrac{\phi}{2}\bigg)\sin\bigg(2\pi x-20t+\dfrac{\phi}{2}\bigg)"


Phase difference = "\phi"

For destructive interference,

Amplitude of resultant wave = 0

"\therefore\cos\dfrac{\phi}{2}=0"

"\phi=n\pi"

for Path difference, "\Delta=\dfrac{\lambda}{2\pi}\phi"

"\lambda=1"

"\Delta=\dfrac{1}{2\pi}n\pi"

for n=1

"\Delta=\dfrac{1}{2}\space m"